A population consists of the four number (2,3,6,9).

Consider all possible samples of size 2 that can be drawn with replacement from this population. Answer the following:


a. List all possible samples of size 2 which can be drawn with replacement from this population.

b. Compute the Population mean.

c. Compute the Population Standard deviation.

d. Find the mean of the sampling distribution.



1
Expert's answer
2022-05-06T03:05:26-0400

a. The number of possible samples which can be selected with replacement is

"N^n=4^2=16."

All the possible samples of sizes n=2 wich can be drawn with replacement from the population:

"\{ (2,2), (2,3),(2,6),(2,9),\\
(3,2),(3,3),(3,6),(3,9),\\
(6,2), (6,3),(6,6),(6,9),\\
(9,2),(9,3),(9,6),(9,9)\}."


b. The population mean:

"\mu=\cfrac{2+3+6+9}{4}=5."


c. The population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 2-5,3-5,6-5,9-5
\end{Bmatrix}="

"=\begin{Bmatrix}
-3, - 2,1,4
\end{Bmatrix},"

"\sigma^2=(-3)^2\cdot \cfrac{1}{4}+(-2)^2\cdot \cfrac{1}{4}+\\
+1^2\cdot \cfrac{1}{4}+4^2\cdot \cfrac{1}{4}=7.5."

The population standard deviation:

"\sigma=\sqrt{7.5}=2.739."


d. The mean of the sampling distribution of sample means:

"\mu_{\bar x} =\mu=5."